# Variational inference via radial transport

### v1

@@ Abstract
In variational inference (VI), the practitioner approximates a high-dimensional distribution $\pi$ with a simple surrogate one, often a (product) Gaussian distribution. However, in many cases of practical interest, Gaussian distributions might not capture the correct radial profile of $\pi$, resulting in poor coverage. In this work, we approach the VI problem from the perspective of optimizing over these radial profiles. Our algorithm radVI is a cheap, effective add-on to many existing VI schemes, such as Gaussian (mean-field) VI and Laplace approximation. We provide theoretical convergence guarantees for our algorithm, owing to recent developments in optimization over the Wasserstein space—the space of probability distributions endowed with the Wasserstein distance—and new regularity properties of radial transport maps in the style of Caffarelli (2000).
@@ 1 Introduction
Variational inference (VI) is a fundamental optimization problem that takes place over subsets of probability distributions (Wainwright and Jordan, 2008; Blei et al., 2017). We consider a standard setup that arises in many applications, where the practitioner is given a high-dimensional posterior distribution $\pi\propto\exp(-V)$ and the goal is to solve
@@
where $\mathcal{C}\subset\mathcal{P}(\mathbb{R}^{d})$ is a fixed set of probability distributions. VI is a powerful computational stand-in for standard Markov Chain Monte Carlo (MCMC) methods for sampling from unnormalized posteriors $\pi$. Indeed, while MCMC methods require simulating Markov chains for prohibitively long periods of time, it might be possible to instead quickly learn a surrogate density that is a good enough approximation to the posterior for practical purposes; see the review by Blei et al. (2017) for more details.
In VI, the choice of $\mathcal{C}\subset\mathcal{P}(\mathbb{R}^{d})$ is of the utmost importance. For example, the case where $\mathcal{C}$ is the set of all Gaussians (with positive definite covariance) is known as Gaussian VI (Barber and Bishop, 1997; Seeger, 1999; Opper and Archambeau, 2009). In large-scale machine learning applications, it is also common to optimize over the class of Gaussians with diagonal covariance, resulting in mean-field Gaussian VI. While these algorithms have a long history, a rigorous, theoretical analysis is only just emerging, based on the theory of optimal transport through Wasserstein gradient flows (Ambrosio et al., 2008). For example, the Gaussian case has been studied by Lambert et al. (2022); Diao et al. (2023); Kim et al. (2024). We note that it is possible to implement algorithms based on mixtures of Gaussians as outlined by Lambert et al. (2022); Petit-Talamon et al. (2025), though the mathematical analysis in this case is significantly more challenging. Separately, Laplace approximation is an alternative means of obtaining a surrogate measure to $\pi$, where one considers the following Gaussian approximation $\mathcal{N}(x^{\star},(\nabla^{2}V(x^{\star}))^{-1})$ where $x^{\star}=\text{argmin}\,\,V$. The literature on Laplace approximations is vast; see e.g., Robert and Casella (2004).
Margossian and Saul (2025) highlight the strengths and weaknesses of existing VI-based algorithms. Notably, they provide some characterizations for when VI can hope to exactly recover the mean and correlation matrix of a target distribution $\pi$. While they are only interested in these particular statistics, a key detail of the paper is that the variational approximating family must be decided in advance, which leads to demonstrable shortcomings even in small-scale examples.
@@ 1.1 Contributions
as $h$ ranges over non-negative functions on $0,\infty)$. If $m$ and $\Sigma$ are known, or if estimates thereof can be imputed, then we can assume $m=0$ and $\Sigma=I$ via a whitening procedure (see Section 4.5). We henceforth assume that this has been done, so that our variational family is the set $\mathcal{C}_{\rm rad}$ of radially symmetric distributions. This family encompasses the standard Gaussian via $h(y)=\exp(-y/2)$, but also Student-$t$ distributions, the non-smooth Laplace distribution, the logistic distribution, among others.***On the other hand, if we fix $h$ and vary $(m,\Sigma)$, we end up at the theory of elliptical families, see e.g., Muirhead (1982, Section 1.4).
@@
where $\pi\propto\exp(-V)$. Our contributions are of both theoretical and computational interest. We stress that our only assumptions throughout this work will be on the true posterior $\pi$, namely that $\pi$ is log-smooth and strongly log-concave and centered at the origin. This pair of assumptions has been leveraged in nearly all works on the theory of sampling (Chewi, 2026) and in the theoretical and computational study of variational inference (Lambert et al., 2022; Arnese and Lacker, 2024; Lacker et al., 2024; Lavenant and Zanella, 2024; Jiang et al., 2025).
In Section 3, we prove existence and uniqueness of the radial minimizer $\pi_{\rm rad}^{\star}$, as well as establish regularity properties of said minimizer. For example, if $\pi$ is log-smooth and strongly log-concave, Theorem 3.4 states that $\pi_{\rm rad}^{\star}$ is as well. We also prove Caffarelli-type contraction estimates (Caffarelli, 2000) for the corresponding optimal radial transport map $T_{\rm rad}^{\star}$ from the standard Gaussian $\rho=\mathcal{N}(0,I)$, say, to $\pi_{\rm rad}^{\star}$; see Theorem 3.5.
@@ 2.1 Primer on optimal transport
where $\Pi(\mu,\nu)$ is the set of joint measures with first marginal $\mu$ and second marginal equal to $\nu$. For more details, see Villani (2009); Santambrogio (2015); Chewi et al. (2025).
@@
where $\mathcal{T}(\mu,\nu)$ is the set of admissible transport maps between $\mu$ and $\nu$, which consist of vector-valued functions $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that for $X\sim\mu$, it holds that $T(X)\sim\nu$. Without further assumptions, it is possible that such a $T^{\star}$ need not exist. A theorem due to Brenier (1991) unifies both (1) and (2).
@@ 2.2 Geodesic convexity via compatible families of optimal transport maps
where $\Psi:[0,\infty)\to[0,\infty)$ is strictly monotone and continuous. One can check that $\mathcal{T}_{\rm rad}$ forms a compatible family of optimal transport maps (see Panaretos and Zemel, 2020, Section 2.3.2), and thus $(\mathcal{T}_{\rm rad})_{\sharp}\rho$ forms a geodesically convex subset of the Wasserstein space.
For an introduction to the space $\mathbb{W}_{2}$ focused on applications to statistics, see Panaretos and Zemel (2020); Chewi et al. (2025).
@@ 3 Setup and theoretical results
and minimized at the origin. These conditions are standard in the theoretical literature on log-concave sampling (Chewi, 2026) as well as in variational inference (see, e.g., Lambert et al., 2022; Arnese and Lacker, 2024; Lacker et al., 2024; Lavenant and Zanella, 2024; Jiang et al., 2025).
@@ 3.2 Regularity of optimal radial maps
We now establish regularity properties of optimal radial (transport) maps from, say, the standard Gaussian to the minimizer of ($\mathsf{radVI}$). Our main theorem below is essentially a specialization of Caffarelli’s contraction theorem (Caffarelli, 2000) to the radial component of the optimal transport map.
@@ 4 Parametrize then optimize
We follow the approach put forth by Jiang et al. (2025), who suggest to appropriately parametrize the space of transport maps and then to optimize over it. Letting $\{\Psi_{j}\}_{j=0}^{J}$ be a fixed set of basis functions and $\alpha>0$, we consider a parametrized subset of transport maps $\mathcal{T}_{J}\subseteq\mathcal{T}_{\rm rad}$, given by
@@ 4.1 Approximation guarantees
Our first result of this section establishes a “universality” property of the set $\mathcal{T}_{J}$ from an approximation perspective, and yields the choice of $R$ and $J$ required to complete our construction.
@@ 4.2 Proposed algorithm: radVI
It is worth stressing that the choice of using gradient descent as a first-order algorithm was entirely arbitrary, and many other algorithms (e.g., Frank–Wolfe, mirror descent, etc.) are applicable within our framework. The functional of interest can be suitably arbitrary as well; see Jiang et al. (2025) for more details.
@@ 4.3 Optimization guarantees
See Jiang et al. (2025, Appendix C) for a proof of this fact (in particular, the proof of their Theorem 5.9 and Corollary C.3). Thus, for $J$ large enough, the minimizer over $\mathcal{T}_{J}$, i.e., $T_{\lambda^{\star}}$, can be made arbitrarily close to $T^{\star}_{\rm rad}$. To conclude, it remains to quantitatively assess that $T_{\lambda^{(k)}}\to T_{\lambda^{\star}}$, and then string everything together.
@@
where we recall that $Q\succ 0$ is the fixed Gram matrix defined from the basis functions. If $\lambda\mapsto\mathcal{F}(T_{\lambda})$ is smooth and strongly convex (with respect to $\|\cdot\|_{Q}$), we can readily apply existing results for the convergence of first-order algorithms (see, e.g., Beck, 2017). Thankfully, the requisite properties of our functional can be verified; see Appendix C.3.
@@ 4.4 Stochastic optimization
Note that this method fails if $\nabla^{2}V(x^{\star})$ is not invertible, and is known to be inaccurate when $\pi$ is heavily skewed (when the mode is far away from the mean). See Katsevich (2023, 2024) for recent statistical developments.
@@
see Lambert et al. (2022); Diao et al. (2023); Kim et al. (2023). A major limitation to GVI is the storage and per-iteration complexity, as the running covariance estimate needs to be stored and inverted at each iteration, which is costly for $d\gg 1$. Optimizing over product Gaussian measures (Gaussian mean-field VI, or MFVI) can mitigate these numerical hurdles, reducing the per-iteration complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d)$ at the cost of possibly being much farther from $\pi$.
@@
We briefly mention the work of Liang et al. (2022), which similarly tries to adjust the tail distributions of their approximated distribution. However, their approach is (i) parametric in nature, and (ii) builds off the mean-field variational inference perspective. Specifically, they only consider product measure approximations to the posteriors.
@@ 5 Experiments
In all experiments, $d=50$, we choose $\alpha=0.01$ and $\lambda^{(0)}=\bm{1}_{J}$, $R=\sqrt{\log d}$, and $\delta=d^{-1/6}$ as parameters for constructing our dictionary. For LA, we use a standard optimization solver and closed-form expressions of the gradient and Hessian. For GVI, we use the Forward–Backward method of Diao et al. (2023).The code used for our implementation of radVI and the numerical experiments are made available at github.com/gluca99/radVI.
@@
Table 1: Estimated Wasserstein distance between various VI solutions for learning isotropic targets.
@@ 5.1 Learning isotropic radial families
Finally, we mention that we additionally performed a simple sensitivity analysis regarding our hyperparameter $\alpha$ similar to Figure 2 of Jiang et al. (2025). Focusing on the isotropic Gaussian case in which the ground-truth value is $\alpha=1/\sqrt{L_{V}}=1$, we run our algorithm where we vary $\alpha\in\{10^{0},10^{-1},10^{-2}\}$. Figure 3 shows that radVI converges to the optimal parameters exponentially fast, highlighting how our algorithm is robust to the choice of $\alpha$.
@@ 5.2 Learning anisotropic distributions
To investigate how radVI can be used to improve GVI for parameter estimation, we study Neal’s funnel distribution, a common example in the VI literature (Neal, 2003; Betancourt and Girolami, 2015). For $d>1$, suppose $z\sim\mathcal{N}(0,4)$ and $x_{i}\sim\mathcal{N}(0,e^{z})$ for $i\in[d]$; one can explicitly write $\pi\in\mathcal{P}(\mathbb{R}^{d+1})$ and compute the corresponding log-density and its derivatives. We stress that this model is misspecified for radial distributions, and instead fall under the category of a structured posterior (Sheng et al., 2025). We follow related work and estimate the following quantities from the posterior: $\mathbb{E}[z^{2}]=4$, $\mathbb{E}[x_{1}^{2}]\approx 7.389$, and $\mathbb{P}(z>2)\approx 0.51$.
@@
Parameter $\mathbb{E}[z^{2}]=4$ $\mathbb{E}[x_{1}^{2}]\approx 7.389$ $\mathbb{P}(z>2)\approx 0.51$
@@
Parameter $\mathbb{E}[z^{2}]=4$ $\mathbb{E}[x_{1}^{2}]\approx 7.389$ $\mathbb{P}(z>2)\approx 0.51$
@@
Table 2: In $d\in\{25,50\}$, we compare the performance of GVI and GVI+radVI for parameter estimation; in parentheses we include the effective sample size.

### v2

@@ Abstract
In variational inference (VI), the practitioner approximates a high-dimensional distribution $\pi$ with a simple surrogate one, often a (product) Gaussian distribution. However, in many cases of practical interest, Gaussian distributions might not capture the correct radial profile of $\pi$, resulting in poor coverage. In this work, we approach the VI problem from the perspective of optimizing over these radial profiles. Our algorithm radVI is a cheap, effective add-on to many existing VI schemes, such as Gaussian (mean-field) VI and Laplace approximation. We provide theoretical convergence guarantees for our algorithm, owing to recent developments in optimization over the Wasserstein space—the space of probability distributions endowed with the Wasserstein distance—and new regularity properties of radial transport maps in the style of Caffarelli, (2000).
@@ 1 Introduction
Variational inference (VI) is a fundamental optimization problem that takes place over subsets of probability distributions (Wainwright and Jordan,, 2008; Blei et al.,, 2017). We consider a standard setup that arises in many applications, where the practitioner is given a high-dimensional posterior distribution $\pi\propto\exp(-V)$ and the goal is to solve
@@
where $\mathcal{C}\subset\mathcal{P}(\mathbb{R}^{d})$ is a fixed set of probability distributions. VI is a powerful computational stand-in for standard Markov Chain Monte Carlo (MCMC) methods for sampling from unnormalized posteriors $\pi$. Indeed, while MCMC methods require simulating Markov chains for prohibitively long periods of time, it might be possible to instead quickly learn a surrogate density that is a good enough approximation to the posterior for practical purposes; see the review by Blei et al., (2017) for more details.
In VI, the choice of $\mathcal{C}\subset\mathcal{P}(\mathbb{R}^{d})$ is of the utmost importance. For example, the case where $\mathcal{C}$ is the set of all Gaussians (with positive definite covariance) is known as Gaussian VI (Barber and Bishop,, 1997; Seeger,, 1999; Opper and Archambeau,, 2009). In large-scale machine learning applications, it is also common to optimize over the class of Gaussians with diagonal covariance, resulting in mean-field Gaussian VI. While these algorithms have a long history, a rigorous, theoretical analysis is only just emerging, based on the theory of optimal transport through Wasserstein gradient flows (Ambrosio et al.,, 2008). For example, the Gaussian case has been studied by Lambert et al., (2022); Diao et al., (2023); Kim et al., (2023). We note that it is possible to implement algorithms based on mixtures of Gaussians as outlined by Lambert et al., (2022); Petit-Talamon et al., (2025), though the mathematical analysis in this case is significantly more challenging. Separately, Laplace approximation is an alternative means of obtaining a surrogate measure to $\pi$, where one considers the following Gaussian approximation $\mathcal{N}(x^{\star},(\nabla^{2}V(x^{\star}))^{-1})$ where $x^{\star}=\text{argmin}\,\,V$. The literature on Laplace approximations is vast; see e.g., Robert and Casella, (2004).
Margossian and Saul, (2025) highlight the strengths and weaknesses of existing VI-based algorithms. Notably, they provide some characterizations for when VI can hope to exactly recover the mean and correlation matrix of a target distribution $\pi$. While they are only interested in these particular statistics, a key detail of the paper is that the variational approximating family must be decided in advance, which leads to demonstrable shortcomings even in small-scale examples.
@@ 1.1 Contributions
as $h$ ranges over non-negative functions on $0,\infty)$. If $m$ and $\Sigma$ are known, or if estimates thereof can be imputed, then we can assume $m=0$ and $\Sigma=I$ via a whitening procedure (see Section 4.5). We henceforth assume that this has been done, so that our variational family is the set $\mathcal{C}_{\rm rad}$ of radially symmetric distributions. This family encompasses the standard Gaussian via $h(y)=\exp(-y/2)$, but also Student-$t$ distributions, the non-smooth Laplace distribution, the logistic distribution, among others.***On the other hand, if we fix $h$ and vary $(m,\Sigma)$, we end up at the theory of elliptical families, see e.g., Muirhead, (1982, Section 1.4).
@@
where $\pi\propto\exp(-V)$. Our contributions are of both theoretical and computational interest. We stress that our only assumptions throughout this work will be on the true posterior $\pi$, namely that $\pi$ is log-smooth and strongly log-concave and centered at the origin. This pair of assumptions has been leveraged in nearly all works on the theory of sampling (Chewi,, 2026) and in the theoretical and computational study of variational inference (Lambert et al.,, 2022; Arnese and Lacker,, 2024; Lacker et al.,, 2024; Lavenant and Zanella,, 2024; Jiang et al.,, 2025).
In Section 3, we prove existence and uniqueness of the radial minimizer $\pi_{\rm rad}^{\star}$, as well as establish regularity properties of said minimizer. For example, if $\pi$ is log-smooth and strongly log-concave, Proposition 3.4 states that $\pi_{\rm rad}^{\star}$ is as well. We also prove Caffarelli-type contraction estimates (Caffarelli,, 2000) for the corresponding optimal radial transport map $T_{\rm rad}^{\star}$ from the standard Gaussian $\rho=\mathcal{N}(0,I)$, say, to $\pi_{\rm rad}^{\star}$; see Theorem 3.5.
@@ 2.1 Primer on optimal transport
where $\Pi(\mu,\nu)$ is the set of joint measures with first marginal $\mu$ and second marginal equal to $\nu$. For more details, see Villani, (2009); Santambrogio, (2015); Chewi et al., (2025).
@@
where $\mathcal{T}(\mu,\nu)$ is the set of admissible transport maps between $\mu$ and $\nu$, which consist of vector-valued functions $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that for $X\sim\mu$, it holds that $T(X)\sim\nu$. Without further assumptions, it is possible that such a $T^{\star}$ need not exist. A theorem due to Brenier, (1991) unifies both (1) and (2).
@@ 2.2 Geodesic convexity via compatible families of optimal transport maps
where $\Psi:[0,\infty)\to[0,\infty)$ is strictly monotone and continuous. One can check that $\mathcal{T}_{\rm rad}$ forms a compatible family of optimal transport maps (see Panaretos and Zemel,, 2020, Section 2.3.2), and thus $(\mathcal{T}_{\rm rad})_{\sharp}\rho$ forms a geodesically convex subset of the Wasserstein space.
For an introduction to the space $\mathbb{W}_{2}$ focused on applications to statistics, see Panaretos and Zemel, (2020); Chewi et al., (2025).
@@ 3 Setup and theoretical results
and minimized at the origin. These conditions are standard in the theoretical literature on log-concave sampling (Chewi,, 2026) as well as in variational inference (see, e.g., Lambert et al.,, 2022; Arnese and Lacker,, 2024; Lacker et al.,, 2024; Lavenant and Zanella,, 2024; Jiang et al.,, 2025).
@@ 3.2 Regularity of optimal radial maps
We now establish regularity properties of optimal radial (transport) maps from, say, the standard Gaussian to the minimizer of ($\mathsf{radVI}$). Our main theorem below is essentially a specialization of Caffarelli’s contraction theorem (Caffarelli,, 2000) to the radial component of the optimal transport map.
@@ 4 Parametrize then optimize
We follow the approach put forth by Jiang et al., (2025), who suggest to appropriately parametrize the space of transport maps and then to optimize over it. Letting $\{\Psi_{j}\}_{j=0}^{J}$ be a fixed set of basis functions and $\alpha>0$, we consider a parametrized subset of transport maps $\mathcal{T}_{J}\subseteq\mathcal{T}_{\rm rad}$, given by
@@ 4.1 Approximation guarantees
Our first result of this section establishes a “universality” property of the set $\mathcal{T}_{J}$ from an approximation perspective, and yields the choice of $R$ and $J$ required to complete our construction. Since we mainly care about high-dimensional approximation, henceforth we assume $d\geq 3$.
@@ 4.2 Proposed algorithm: radVI
It is worth stressing that the choice of using gradient descent as a first-order algorithm was entirely arbitrary, and many other algorithms (e.g., Frank–Wolfe, mirror descent, etc.) are applicable within our framework. The functional of interest can be suitably arbitrary as well; see Jiang et al., (2025) for more details.
@@ 4.3 Optimization guarantees
See Jiang et al., (2025, Appendix C) for a proof of this fact (in particular, the proof of their Theorem 5.9 and Corollary C.3). Thus, for $J$ large enough, the minimizer over $\mathcal{T}_{J}$, i.e., $T_{\lambda^{\star}}$, can be made arbitrarily close to $T^{\star}_{\rm rad}$. To conclude, it remains to quantitatively assess that $T_{\lambda^{(k)}}\to T_{\lambda^{\star}}$, and then string everything together.
@@
where we recall that $Q\succ 0$ is the fixed Gram matrix defined from the basis functions. If $\lambda\mapsto\mathcal{F}(T_{\lambda})$ is smooth and strongly convex (with respect to $\|\cdot\|_{Q}$), we can readily apply existing results for the convergence of first-order algorithms (see, e.g., Beck,, 2017). Thankfully, the requisite properties of our functional can be verified; see Appendix C.3.
@@ 4.4 Stochastic optimization
Note that this method fails if $\nabla^{2}V(x^{\star})$ is not invertible, and is known to be inaccurate when $\pi$ is heavily skewed (when the mode is far away from the mean). See Katsevich, (2023, 2024) for recent statistical developments.
@@
see Lambert et al., (2022); Diao et al., (2023); Kim et al., (2023). A major limitation to GVI is the storage and per-iteration complexity, as the running covariance estimate needs to be stored and inverted at each iteration, which is costly for $d\gg 1$. Optimizing over product Gaussian measures (Gaussian mean-field VI, or MFVI) can mitigate these numerical hurdles, reducing the per-iteration complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d)$ at the cost of possibly being much farther from $\pi$.
@@
We briefly mention the work of Liang et al., (2022), which similarly tries to adjust the tail distributions of their approximated distribution. However, their approach is (i) parametric in nature, and (ii) builds off the mean-field variational inference perspective. Specifically, they only consider product measure approximations to the posteriors.
@@ 5 Experiments
In all experiments, $d=50$, we choose $\alpha=0.01$ and $\lambda^{(0)}=\bm{1}_{J}$, $R=\sqrt{\log d}$, and $\delta=d^{-1/6}$ as parameters for constructing our dictionary. For LA, we use a standard optimization solver and closed-form expressions of the gradient and Hessian. For GVI, we use the Forward–Backward method of Diao et al., (2023).The code used for our implementation of radVI and the numerical experiments are made available at github.com/gluca99/radVI.
@@
Table 1: Estimated squared Wasserstein distance between various VI solutions for learning isotropic targets.
@@ 5.1 Learning isotropic radial families
Finally, we mention that we additionally performed a simple sensitivity analysis regarding our hyperparameter $\alpha$ similar to Figure 2 of Jiang et al., (2025). Focusing on the isotropic Gaussian case in which the ground-truth value is $\alpha=1/\sqrt{L_{V}}=1$, we run our algorithm where we vary $\alpha\in\{10^{0},10^{-1},10^{-2}\}$. Figure 3 shows that radVI converges to the optimal parameters exponentially fast, highlighting how our algorithm is robust to the choice of $\alpha$.
@@ 5.2 Learning anisotropic distributions
To investigate how radVI can be used to improve GVI for parameter estimation, we study Neal’s funnel distribution, a common example in the VI literature (Neal,, 2003; Betancourt and Girolami,, 2015). For $d>1$, suppose $z\sim\mathcal{N}(0,4)$ and $x_{i}\sim\mathcal{N}(0,e^{z})$ for $i\in[d]$; one can explicitly write $\pi\in\mathcal{P}(\mathbb{R}^{d+1})$ and compute the corresponding log-density and its derivatives. We stress that this model is misspecified for radial distributions, and instead fall under the category of a structured posterior (Sheng et al.,, 2025). We follow related work and estimate the following quantities from the posterior: $\mathbb{E}[z^{2}]=4$, $\mathbb{E}[x_{1}^{2}]\approx 7.389$, and $\mathbb{P}(|z|>2)\approx 0.317$.
@@
Parameter $\mathbb{E}[z^{2}]=4$ $\mathbb{E}[x_{1}^{2}]\approx 7.389$ $\mathbb{P}(|z|>2)\approx 0.317$
@@
Parameter $\mathbb{E}[z^{2}]=4$ $\mathbb{E}[x_{1}^{2}]\approx 7.389$ $\mathbb{P}(|z|>2)\approx 0.317$
@@
Table 2: In $d\in\{25,50\}$, we compare the performance of GVI and GVI+radVI for parameter estimation.